OpenStudy (anonymous):

Are the lines parallel, perpendicular, or neither? 5x + 2y = 10 15x + 4y = -4

Nnesha (nnesha):

hmm hey :=)) i was about to leave.. well hmm rewrite both equation in slope intercept form y=mx+B if the slopes are the same then parallel (parallel lines have same slope ) if the slopes are negative reciprocal ( exxample: a/b and the other slope -b/a) then perpendicular

Nnesha (nnesha):

can you solve for y ?? :=)) 5x+2y=10 btw y=mx+b m is slope

OpenStudy (anonymous):

Refer to the attachment that includes a plot.